# Machine Unlearning: A Comprehensive Survey

### v2

@@ 6.2. The Employed Datasets
7. Privacy and Security Threats on Machine Unlearning
@@
Although unlearning methods were initially proposed to safeguard users’ privacy, many researchers have noticed that unlearning brings new privacy threats simultaneously. Similar to machine learning scenarios, there are two main kinds of privacy threats in unlearning. They are membership inference attacks and private information reconstruction attacks. In this section, we detail the two prevalent unlearning privacy threats.
@@ 7.1.2. Privacy Reconstruction Attacks in Unlearning
Zanella-Béguelin et al. (Zanella-Béguelin et al., 2020) indicated that the releasing snapshot of overlapped language models would leak the privacy of the training dataset. They verified that the model updates significantly threaten the private information added to or deleted from the training dataset by many experimental results. Zanella-Béguelin et al. found five phenomenons. First, an attacker can extract particular sentences used or not in the training dataset by comparing two models. Second, analyzing more model snapshots shows more information about the updated data than considering fewer model snapshots. Third, adding or deleting other non-private data during model updates can not mitigate privacy leakage. Fourth, differential privacy can reduce privacy leakage risks and decrease trained models’ accuracy. Fifth, to mitigate the privacy leakage risks while keeping the model utility, the server can limit the model parameters access or only output a subset of the results.
Many studies further utilized these privacy threats to evaluate the unlearning effect. Huang et al. (Huang et al., 2021a) proposed Ensembled Membership Auditing (EMA) for auditing data erasure. They use the membership inference to assess the removing effectiveness of unlearning. Graves et al. (Graves et al., 2021) indicated that if an attacker can infer the sensitive information that was wanted to be erased, then it means that the server has not guarded the rights to be forgotten. Baumhauer et al. (Baumhauer et al., 2022) developed linear filtration to sanitize classification models with logits prediction after class-wide deletion requests. They verified their methods by testing how well the method defends against privacy attacks.
Both (Chen et al., 2021) and (Zanella-Béguelin et al., 2020) pointed out that differential privacy guarantees that a model does not reveal too much knowledge about any training sample. A DP-protected model can further guarantee the group’s privacy by binding the impacts of a bunch of training samples. If using DP to protect the privacy of a bunch of $|D\backslash D_{e}|$training examples against snapshot attacks on $\theta_{D},\theta_{D\backslash D_{e}}$, it means that $\theta_{D\backslash D_{e}}$ cannot be more useful than $\theta_{D}$.
@@ 7.2. Security Threats on Machine Unlearning
There are several papers that address security threats targeting machine unlearning. For instance, Marchant et al. (Marchant et al., 2022) examine an attacker’s strategy to increase the computational cost of performing machine unlearning. Studies by Zhao et al. (Zhao et al., 2024) and Hu et al. (Hu et al., ) investigate security vulnerabilities within unlearning systems, specifically by uploading customized malicious data update requests to negatively influence the model utility. Additionally, some studies propose backdooring or poisoning methods that exploit unlearning requests to achieve backdoor insertion, as discussed by (Di et al., 2022) and (Liu et al., 2024). These works collectively highlight the need for robust security measures in machine unlearning to mitigate these emerging threats.
@@ 8. Machine Unlearning Applications
(Wang et al., 2019) Mitigate backdoor - Median Absolute Deviation L1-Norm, ##\##FP, ACC, ASR 2019
(Du et al., 2019) Anomaly detection - Maintain memory set ##\##FP and ##\##FN 2019
(Cao et al., 2018) Repair pollution KARMA Cluster and unlearn DA 2018
@@
  • ##\##FN: number of false negative, ##\##FP: number of false positive
@@
Study (Wang et al., 2019) applied machine unlearning in detecting and mitigating backdoor attacks in DNNs. They designed two approaches to eliminate backdoored neurons from the backdoored model and repair the model to be strong against malicious pictures. Then, in (Liu et al., 2022b), they offered that a backdoor model learned the poisoned decision boundary. Data points with triggers are all classified into the target class. They reversed the backdoor injection process to defend against it in machine unlearning, which is simple but effective. Their method contains two primary steps: first, they utilize a max-entropy staircase approximator to complete trigger reconstruction; second, they remove the added triggers using unlearning. They named these two key steps of BAERASER as trigger pattern recovery and trigger pattern unlearning. Via a dynamic penalty mechanism, they mitigated the sharp accuracy degradation of gradient-ascent-based machine unlearning methods.
Repairing pollution is another successful unlearning application. Cao et al. (Cao et al., 2018) proposed “KARMA” to search various subsets of original datasets and return the subset with the highest misclassifications. First, KARMA searches for possible reasons that lead to the wrong ML model classification. It clusters the misclassified samples into various domains and extracts the middle of clusters. KARMA prioritizes the search for matching examples in the original datasets using these extracted centers. Second, KARMA grows the reason discovered in the first step by discovering more training samples and creating a cluster. Third, KARMA determines if a causality cluster is polluted and calculates how many samples the cluster contains.
Machine unlearning was also applied to lifelong anomaly detection problems. Du et al. (Du et al., 2019) proposed new approaches to solve the challenges of exploding loss and sharp accuracy degradation caused by unlearning. In their presentation, they explain their techniques, emphasizing handling false negatives. When meeting such case $x_{t}$, their purpose is, therefore, to modify the model to reduce the possibility of $p(x_{t})$, or to increase the loss $\mathcal{L}(x_{t})$.
Huang et al. (Huang et al., 2021b) presented a method that can make samples unlearnable by injecting error-minimizing noise. This noise is intentionally synthesized to diminish the error of the samples close to zero, which can mislead the model into considering there is ”nothing” to learn. They first tried the error-maximizing noise but found that it could not prevent DNN learning when used sample-wise to the training data points. Therefore, they then begin to study the opposite direction of error-maximizing noise. In particular, they proposed the error-minimizing noise to stop the model from being punished by the loss function during traditional ML model training. Therefore, it can mislead the ML model to consider that there is ”nothing” to learn.
@@ 9. Lessons Learnt and Discussions
In the research domain of machine unlearning, researchers mainly face three difficult challenges, the stochasticity of training, incrementality of training, and catastrophe of unlearning, when they try to design efficient and effective unlearning mechanisms replacing naive retraining. They tried many mechanisms to mitigate the influence of these challenges. For example, in exact unlearning, they designed split algorithms that divide the final model into several sub-models and avoid the stochasticity and incrementality of one sub-model to influence other sub-models. In approximate unlearning, they bounded the removed estimation to avoid accuracy degradation. However, it is still not easy to analyze them clearly and solve them totally. Moreover, as the research in machine unlearning in-depth, researchers extended it to new situations, such as federated learning and graph learning, and they met corresponding new challenges. Here, we discuss the differences between various unlearning scenarios and their corresponding challenges. Based on these challenges, we list some potential research directions on which we can focus.
In the common centralized unlearning scenario, retraining from scratch can achieve the best unlearning effect, but it is expensive in both computation and storage. Existing methods try to design new unlearning mechanisms to reduce the cost from the two aspects. Although they proposed many methods, they just mitigated the influence of the main challenges and still have not solved them. Existing unlearning methods were at the beginning phase, which is trying to implement effective unlearning. They proposed mechanisms, but most of them cannot guarantee the final results, so they also bound the unlearned item (Guo et al., 2020) or set an unlearning threshold (Nguyen et al., 2020). Although they also introduced many metrics to verify the unlearning effect, the unlearning and verification processes are split. This means that the unlearning result is uncertain during the unlearning period before the verification is finished. Therefore, we find and list some open questions in centralized unlearning.
@@
    Optimizing the existing challenges. The first two challenges are hard to solve exactly, so recent research focuses on reducing utility degradation after unlearning. Even though many studies exist, there is still a long way to go to eliminate the degradation of unlearning.
@@
    How do we unlearn with a certain or exact goal? In other words, can we unlearn as learning? If we know the purpose, to what extent we have unlearned, and when we can finish or stop the unlearning process.
@@
    Federated unlearning cannot use the fast retraining methods because data is out of reach for the server. Therefore, federated unlearning can only be implemented using approximate unlearning methods. However, as we know, approximate unlearning easily causes catastrophic unlearning, and federated learning is more vulnerable to degradation, so controlling the catastrophic in federated unlearning will be more urgent than in centralized unlearning.
@@
    Existing federated unlearning methods require the activation of all users, including those without unlearning requests, to assist in the unlearning process. These approaches are impractical and inefficient, particularly when unlearning requests are frequent. Therefore, it is crucial to study how to balance the unlearning effect and efficiency in federated unlearning.

### v3

@@ 5.2. The Employed Datasets
6. Domain-Centric Machine Unlearning
After introducing the classic centralized machine unlearning techniques and corresponding evaluation methods, in this section, we will introduce the studies of domain-centric unlearning, such as federated unlearning, graph unlearning, diffusion model unlearning, and large language model unlearning.
6.1. Federated Unlearning
FL was initially introduced to protect the privacy of participating clients during the machine learning training process in distributed settings. All participants will only upload their locally trained model parameters instead of their sensitive local data to the FL server during model training processes (McMahan et al., 2016). Therefore, in a federated learning scenario, limited access to the dataset will become a unique challenge when implementing unlearning. According to the unlearning target of a whole client’s contribution or samples’ contribution, we can roughly divide existing unlearning studies into two categories: client-level and sample-level federated unlearning. Since the client-level unlearning is usually operated in the server side and the sample-level unlearning usually needs the clients’ participation, we can also call them server-side and client-side federated unlearning, as shown in Figure 7.
Client-level (Server-side) Federated Unlearning. Since the local data cannot be uploaded to the federated learning (FL) server side, most federated unlearning methods try to erase a certain client’s contribution from the trained model by storing and estimating the contribution of uploaded parameters. In this situation, they can implement federated unlearning without interacting with the client, shown as the server-side federated unlearning in Fig. 7 (a). The two representative methods are (Liu et al., 2021; Wu et al., 2022b). Liu et al. (Liu et al., 2021) proposed “FedEraser” to sanitize the impact of a FL client on the global FL model. In particular, during FL training process, the FL-Server maintains the updates of the clients at each routine iteration and the index of the related round to calibrate the retrained updates. Based on these operations, they reconstructed the unlearned FL model instead of retraining a new model from scratch. However, FedEraser can only unlearn one client’s data, which means it must unlearn all the contributions of this specific client’s data. It is unsuitable for a client who wants to unlearn a small piece of his data. Study (Wu et al., 2022b) tried to erase a client’s influence from the FL model by removing the historical updates from the global model. They implemented federated unlearning by using knowledge distillation to restore the contribution of clients’ models, which does not need to rely on clients’ participation and any data restriction.
Smaple-level (Client-side) Federated Unlearning. Different from unlearning a whole client’s influence and unlearning a class, Liu et al. (Liu et al., 2022a; Wang et al., 2023c) investigated how to unlearn data samples in FL, shown as the client-side federated unlearning in Fig. 7 (b). In (Liu et al., 2022a), they first defined a federated unlearning problem and proposed a fast retraining method to withdraw the influence of data from the FL model. Then, they proposed an efficient federated unlearning method following the Quasi-Newton methods and the first-order Taylor approximate method (Liu et al., 2022a). They utilized the practical Fisher Information Matrix to model the Hessian matrix at a low cost. Another similar work based on influence function to implement federated unlearning was introduced in (Wang et al., 2024b). Moreover, in (Wang et al., 2023c), the authors implement federated unlearning based on Bayesian inference, and they propose a parameters self-sharing method to reduce the model utility degradation. Although these methods effectively implement federated unlearning, they still need some benign clients to participate in the unlearning training, which limits the realistic deployment.
6.1.1. Graph Unlearning
We introduce graph unlearning as a representative kind of irregular data unlearning. In (Chien et al., 2022a; Chen et al., 2022b; Cong and Mahdavi, 2022; Zhang et al., 2024a), researchers extend regular data machine unlearning to a graph data scenario. Graph structure data are more complex than standard structured data because graph data include not only the feature information of nodes but also the connectivity information of different nodes, shown in Fig. 8. Therefore, Chien et al. (Chien et al., 2022a) proposed node unlearning, edge unlearning, and both node and edge unlearning for simple graph convolutions (SGC). Besides the different information unlearned in a graph learning problem, they found another challenge associated with feature mixing during propagation, which needs to be addressed to establish provable performance guarantees. They gave the theoretical analysis for certified unlearning of GNNs by illustrating the underlying investigation on their generalized PageRank (GPR) extensions and the example of SGC.
6.2. Unlearning Diffusion Models
Exploring unlearning solutions for diffusion models and large language models (LLMs) is popular in recent years. We will first introduce the literature about unlearning methods for diffusion models and then introduce the LLMs unlearning in the next subsection.
Unlike the classic classification unlearning setting, unlearning for generative models must consider about open-ended conditional output distributions. The unlearned model should cease to reproduce specific training examples and protected artistic styles. Commonly, diffusion model unlearning methods can be organized by what is removed. Firstly, data-centric deletion targets the influence of designated training points and aims to approximate retraining on a pruned dataset (Li et al., 2024b; Alberti et al., 2025; Chen et al., 2024). Secondly, concept-centric deletion removes higher-level semantic notions (e.g., nudity or protected styles) by intervening in the model’s internal representations and denoising dynamics (Wu et al., 2025a; Zhang et al., 2024b; Wu et al., 2025b; Cywiński and Deja, 2025).
Data-centric deletion. A line of the diffusion model unlearning treats unlearning as an approximate data removal problem. The goal of these studies is to match the behaviour of a model retrained from scratch based only on the remaining dataset. Li et al. (Li et al., 2024b) studied machine unlearning for image-to-image generative models. They contributed a general unlearning framework for structured translation models such as diffusion models, VQ-GAN, and MAE. They showed on ImageNet-1k and Places-365 that their method can effectively forget target samples with only a small degradation on retrained ones. Alberti et al. (Alberti et al., 2025) focused on data unlearning in diffusion models. They defined retraining without the target data as the gold standard, then proposed SISS (Subtracted Importance Sampled Scores), a method that combines quality preservation and forgetting strength through importance sampling. Experiments on CelebA-HQ, MNIST with T-shirt, and Stable Diffusion show that SISS achieves a strong trade-off between forgetting and generation quality. Chen et al. (Chen et al., 2024) proposed Score Forgetting Distillation (SFD), a data-free unlearning method for diffusion models. Instead of using original training data, SFD adds an unlearning objective into score distillation, encouraging the model to align the scores of unsafe classes.
Concept-centric deletion. A second line of diffusion unlearning studies targets semantic removal rather than individual data points. They usually aim to erase concepts or styles that may be harmful. EraseDiff (Wu et al., 2025a) formulates concept removal as a constrained optimization problem: the model should move away from denoising trajectories associated with the forget set, while staying close to behaviors needed for retained content. AdvUnlearn (Zhang et al., 2024b) shifts the focus to robustness, integrating adversarial training into the unlearning process so that erased concepts cannot be easily recovered through adversarial prompts. Importantly, it shows that editing the text encoder can yield stronger and more transferable robustness than editing the UNet alone. DoCo (Wu et al., 2025b) further targets generalization and utility preservation by aligning the output domains of sensitive and anchor concepts and using gradient surgery to reduce interference with related benign concepts.
In contrast to these parameter-level approaches, SAeUron (Cywiński and Deja, 2025) offers a more interpretable feature-level intervention by training sparse autoencoders on internal activations and blocking the features associated with unwanted concepts. Thereby, it enables transparent and multi-concept unlearning with competitive robustness and generation quality.
Cross-cutting evaluation. For diffusion models, unlearning evaluation should go beyond checking whether a target concept disappears in a few post-edit samples. A convincing evaluation framework must assess at least four properties: forgetting efficacy, utility preservation, robustness, and durability. Zhang et al. (Zhang et al., 2024c) make this point explicit by introducing UnlearnCanvas, a high-resolution benchmark for object and style unlearning that provides a standardised and automated evaluation pipeline with seven quantitative metrics. They also examine harder settings such as adversarial prompts, finer-grained concept removal, and sequential unlearning, showing that immediate deletion performance alone is insufficient for judging success. Ko et al. (Ko et al., 2024) further show that diffusion unlearning can substantially impair text–image alignment, and therefore argue that evaluation must jointly measure forgetting and alignment rather than treating utility as generic sample quality alone. Their post-unlearning alignment framework is motivated precisely by this tension. George et al. (George et al., 2025) extend evaluation to the temporal dimension by showing that supposedly erased concepts can re-emerge after benign fine-tuning.
These studies suggest that diffusion unlearning should be evaluated not only by how much is forgotten immediately after editing, but also by whether the model retains non-target capabilities, resists adversarial recovery, and remains stably unlearned after subsequent updates.
6.3. Unlearning Large Language Models
Unlearning in large language models (LLMs) seeks to remove the concepts and behaviours, such as copyrighted text, personal information, and unsafe responses. Recent studies conceptualize this problem as eliminating undesirable training influence and the capabilities associated with it under two key requirements: locality and utility, meaning that changes should be confined to the targeted knowledge and general performance on unrelated tasks should remain largely intact (Yao et al., 2024a, b; Liu et al., 2025a). From this perspective, existing approaches can be broadly divided into two categories according to where forgetting is implemented. The first is parametric post-hoc unlearning, which directly updates model parameters so that forgetting becomes embedded within the model itself (Yao et al., 2024b, a; Liu et al., 2024a; Huang et al., 2024a; Wang et al., 2025a). The second is non-parametric or suppression-based unlearning, which leaves the underlying model unchanged and instead prevents undesirable outputs through inference-time intervention or control (Liu et al., 2024b).
Parametric post-hoc unlearning. Most LLMs unlearning studies are about post-hoc unlearning directly on pretrained LLMs by updating weights to suppress targeted information. Large Language Model Unlearning (Yao et al., 2024a) presents unlearning primarily as an alignment mechanism, arguing that targeted forgetting can suppress harmful outputs, remove copyrighted content, and reduce hallucinations using only negative examples. Machine Unlearning of Pre-trained Large Language Models (Yao et al., 2024b), in contrast, takes a more systematic and benchmark-driven view, evaluating multiple unlearning strategies on pre-trained models and showing that gradient-based post-hoc unlearning can be highly efficient relative to retraining, especially when forgetting updates are combined with utility-preserving regularisation on retained data.
Building on early gradient-based approaches, subsequent work develops more targeted unlearning mechanisms to better satisfy the locality requirement. Liu et al. (Liu et al., 2024a) propose Selective Knowledge-negation Unlearning (SKU), a two-stage framework that first identifies harmful knowledge and then explicitly negates it, aiming to suppress harmful responses while preserving normal utility. Wang et al. (Wang et al., 2025a) further refine the granularity of forgetting with SeUL, which performs unlearning at the level of sensitive spans instead of whole sequences and introduces dedicated metrics to quantify sensitive-information deletion. These methods reflect a broader shift from coarse unlearning toward more selective and locality-aware interventions, although they differ in whether forgetting is enforced through parameter negation, output-space offsets, or span-level objective design.
Non-parametric unlearning (inference-time control). Unlike parametric unlearning methods that edit model weights, Liu et al. (Liu et al., 2024b) shift forgetting to the inference stage. Their Embedding-Corrupted (ECO) framework treats unlearning as a conditional control problem: the base model remains intact, while a detector identifies prompts associated with forgotten content and an embedding-level intervention steers generation away from that content. In this sense, ECO does not make forgetting intrinsic to the model; rather, it imposes an “unlearned state” on demand for selected inputs. This design is attractive operationally because it is modular, scalable, and easy to deploy across different LLM backbones, especially in settings where direct access to model parameters is unavailable or repeated fine-tuning would be too costly.
Evaluation infrastructure for LLMs unlearning. Across recent work, evaluation in LLM unlearning is becoming both broader and more demanding. Jin et al. (Jin et al., 2024) contribute a dedicated text-only benchmark, RWKU, which evaluates forgetting of real-world entity knowledge on LLaMA3-Instruct (8B) and Phi-3 Mini-4K-Instruct (3.8B) using membership-inference attacks, adversarial probes, and retain tasks drawn from MMLU, BBH, TruthfulQA, TriviaQA, and AlpacaEval. Liu et al. (Liu et al., 2025a) provide a more general evaluation framework, arguing that unlearning should be assessed not only by immediate deletion success but also by scope precision, privacy leakage, robustness, consistency, and effects on causally unrelated capabilities. Extending this agenda beyond text-only models, Li et al. (Li et al., 2024c) propose SIU for multimodal LLMs and introduce MMUBench, built from concepts sampled from MIKE, image collections from web search, and evaluation on LLaVA 7B/13B together with standard multimodal benchmarks such as GQA, VQA-v2, VizWiz, ScienceQA-IMG, TextVQA, POPE, MMBench, and MM-Vet. Taken together, these studies show that unlearning evaluation is moving from coarse deletion scores toward more rigorous, attack-aware, locality-sensitive, and increasingly multimodal assessment.
7. Privacy and Security Issues on Machine Unlearning
@@
Although unlearning methods were initially proposed to safeguard users’ privacy, many researchers have noticed that unlearning brings new privacy threats simultaneously. In this section, we will discuss these privacy and security studies about machine unlearning.
@@ 7.1.2. Privacy Reconstruction Attacks in Unlearning
Zanella-Béguelin et al. (Zanella-Béguelin et al., 2020) indicated that the releasing snapshot of overlapped language models would leak the privacy of the training dataset. They verified that the model updates significantly threaten the private information added to or deleted from the training dataset by many experimental results. Zanella-Béguelin et al. found five phenomena. First, an attacker can extract particular sentences used or not in the training dataset by comparing two models. Second, analyzing more model snapshots shows more information about the updated data than considering fewer model snapshots. Third, adding or deleting other non-private data during model updates can not mitigate privacy leakage. Fourth, differential privacy can reduce privacy leakage risks and decrease trained models’ accuracy. Fifth, to mitigate the privacy leakage risks while keeping the model utility, the server can limit the model parameters access or only output a subset of the results.
Many studies further utilized these privacy threats to evaluate the unlearning effectiveness. Huang et al. (Huang et al., 2021a) proposed Ensembled Membership Auditing (EMA) for auditing data erasure. They use the membership inference to assess the removing effectiveness of unlearning. Graves et al. (Graves et al., 2021) indicated that if an attacker can infer the sensitive information that was wanted to be erased, then it means that the server has not guarded the rights to be forgotten. Baumhauer et al. (Baumhauer et al., 2022) developed linear filtration to sanitize classification models with logits prediction after class-wide deletion requests. They verified their methods by testing how well the method defends against privacy attacks.
[…]
@@ 7.2. Security Threats on Machine Unlearning
There are several papers that introduced security threats targeting machine unlearning. For instance, Marchant et al. (Marchant et al., 2022) examine an attacker’s strategy to increase the computational cost of performing machine unlearning. Studies by Zhao et al. (Zhao et al., 2024) and Hu et al. (Hu et al., ) investigate security vulnerabilities within unlearning systems, specifically by uploading customized malicious data update requests to negatively influence the model utility. Additionally, some studies propose backdooring or poisoning methods that exploit unlearning requests to achieve backdoor insertion, as discussed by (Di et al., 2022) and (Liu et al., 2024c). Huang et al. (Huang et al., 2024b) introduced an unlearning-activated backdoor attack that uses influence-based analysis to select camouflage to trigger samples, so that the backdoor remains stealthy before unlearning but becomes effective after specific deletions have taken place. These works collectively highlight the need for robust security measures in machine unlearning to mitigate these emerging threats.
7.3. Defending Methods
From a defensive standpoint, only a limited number of studies have begun to design privacy-preserving and robust unlearning methods to defend against these emerging threats. Representative examples are BlindU (Wang et al., 2026) and Compressive Representation Forgetting (CRFU) (Wang et al., 2025d). CRFU mitigates reconstruction attacks and related forms of information leakage attacks by imposing an information-bottleneck-inspired compression objective on the representations of data designated for removal (Wang et al., 2025d). In doing so, CRFU aims to suppress the mutual information between latent representations and the erased inputs, while preserving model utility through remembering objectives and a controllable unlearning rate (Wang et al., 2025d). BlindU provides a stricter privacy-preserving unlearning framework compared with CRFU. Privacy preservation of BlindU is achieved by uploading compressed and differential privacy-enabled data as the unlearning requests, and an unlearning method in BlindU for compressed data is designed correspondingly (Wang et al., 2026). More broadly, this line of work signals a move away from viewing unlearning as a simple post hoc algorithmic adjustment, and toward framing it as a joint optimisation problem that simultaneously addresses forgetting, privacy protection, and robustness.
@@ 8. Machine Unlearning Applications
(Wang et al., 2019) Mitigate backdoor - Median Absolute Deviation L1-Norm, #\#FP, ACC, ASR 2019
(Du et al., 2019) Anomaly detection - Maintain memory set #\#FP and #\#FN 2019
(Cao et al., 2018) Repair pollution KARMA Cluster and unlearn DA 2018
@@
  • #\#FN: number of false negative, #\#FP: number of false positive
@@
Mitigating Anomaly. Study (Wang et al., 2019) applied machine unlearning in detecting and mitigating backdoor attacks in DNNs. They designed two approaches to eliminate backdoored neurons from the backdoored model and repair the model to be strong against malicious pictures. Then, in (Liu et al., 2022b), they offered that a backdoor model learned the poisoned decision boundary. Data points with triggers are all classified into the target class. They reversed the backdoor injection process to defend against it in machine unlearning, which is simple but effective. Their method contains two primary steps: first, they utilize a max-entropy staircase approximator to complete trigger reconstruction; second, they remove the added triggers using unlearning. They named these two key steps of BAERASER as trigger pattern recovery and trigger pattern unlearning. Via a dynamic penalty mechanism, they mitigated the sharp accuracy degradation of gradient-ascent-based machine unlearning methods.
Repairing pollution is another successful unlearning application. Cao et al. (Cao et al., 2018) proposed “KARMA” to search various subsets of original datasets and return the subset with the highest misclassifications. First, KARMA searches for possible reasons that lead to the wrong ML model classification. It clusters the misclassified samples into various domains and extracts the middle of clusters. KARMA prioritizes the search for matching examples in the original datasets using these extracted centers. Second, KARMA grows the reason discovered in the first step by discovering more training samples and creating a cluster. Third, KARMA determines if a causality cluster is polluted and calculates how many samples the cluster contains. Du et al. (Du et al., 2019) also explored unlearning in lifelong anomaly detection and tried to mitigating exploding loss and sharp accuracy degradation caused by unlearning.
Data Unlearnable. Huang et al. (Huang et al., 2021b) presented a method that can make samples unlearnable by injecting error-minimizing noise. This noise is intentionally synthesized to diminish the error of the samples close to zero, which can mislead the model into considering there is ”nothing” to learn. They first tried the error-maximizing noise but found that it could not prevent DNN learning when used sample-wise to the training data points. Therefore, they then begin to study the opposite direction of error-maximizing noise. In particular, they proposed the error-minimizing noise to stop the model from being punished by the loss function during traditional ML model training. Therefore, it can mislead the ML model to consider that there is ”nothing” to learn.
@@ 9. Lessons Learnt and Discussions
In the research domain of machine unlearning, researchers mainly face three difficult challenges, the stochasticity of training, incrementality of training, and catastrophe of unlearning. Researchers tried many mechanisms to mitigate the influence of these challenges. For example, in exact unlearning, they designed split algorithms that divide the final model into several sub-models and avoid the stochasticity and incrementality of one sub-model to influence other sub-models (Bourtoule et al., 2021). In approximate unlearning, they bounded the removed estimation to avoid accuracy degradation (Guo et al., 2020). However, it is still not easy to analyze them clearly and solve them totally. As the research in machine unlearning in-depth, researchers extended it to new situations, such as federated learning and graph learning, and they met corresponding new challenges. Here, we discuss the differences between various unlearning scenarios and their corresponding challenges. Based on these challenges, we list some potential research directions on which we can focus.
In the common centralized unlearning scenario, retraining from scratch can achieve the best unlearning effect, but it is expensive in both computation and storage. Existing methods try to design new unlearning mechanisms to reduce the cost from the two aspects. Although researchers proposed many methods, they just solved the challenges partially and cannot guarantee the final unlearning effect. Here, we find and list some open questions in traditional centralized unlearning.
@@
    Challenge: The stochasticity of training and incrementality of training challenges are hard to solve exactly, so recent research mainly focuses on reducing utility degradation after unlearning. Potential Direction: Exploring the unlearning problem with a theoretical guarantee, such as from a differential geometry perspective, will provide the theoretical contribution and explanation to unlearning.
@@
    Challenge: What unlearning mechanisms should be to forget a certain or exact goal in different models, such as in diffusion model or LLMs? Potential Direction: Exploring to what extent we have to unlearn for different unlearning purposes remains under-explored, and how to utilize the learning mechanism to implement unlearning purposes in also an interesting direction.
@@
    Challenge: Federated unlearning is hard to be implemented by using the fast retraining methods because data is out of reach for the server. Most federated unlearning methods are implemented by using approximate unlearning methods. However, as we know, approximate unlearning easily causes catastrophic unlearning, and federated learning is more vulnerable to degradation. Potential Direction: Controlling the catastrophic in federated unlearning will be more urgent than in centralized unlearning.
@@
    Challenge: Existing federated unlearning methods require the activation of all users, including those without unlearning requests, to assist in the unlearning process. These approaches are impractical and inefficient, particularly when unlearning requests are frequent. Potential Direction: It is crucial to study how to balance the unlearning effectiveness and efficiency in federated unlearning.
@@
Unlearning in LLMs and diffusion models introduces additional challenges beyond those in conventional discriminative models. In these generative models, the target to be removed is often distributed across parameters and entangled with related knowledge, so forgetting a specific sample, entity, or concept does not necessarily prevent the model from reproducing semantically related variants (Liu et al., 2025a; Chen et al., 2024). Existing evaluation is also still immature: LLM unlearning is heavily concentrated on a small number of fixed forget-retain benchmarks (Jin et al., 2024), while diffusion unlearning studies have shown that erased concepts may reappear under nearby or adversarial prompts (George et al., 2025). Therefore, future research should focus on defining unlearning targets and scopes more precisely, such as at the entity, concept, and hierarchical knowledge levels, developing more realistic and robust benchmarks, improving continual and black-box unlearning for real deployment settings, and enhancing robustness and interpretability.